In order to set up for a fiar, Guadelupe divides the gymnasium into four sections by putting out two lines of rope that intersect in the middle of the gymnasium. The regions are numbered from one to four in a clockwise manner. If the angle at the ponit of intersection for region 2 has a measure of 69 degrees, what is the angle measure of region 3?
a) 56 b) 111 c) 138 d) 69
plesea show step for step!!!

Answer by jim_thompson5910(35256)

Let x = unknown angle

x + 69 = 180

x + 69 - 69 = 180 - 69

x = 111

So the angle is 111 degrees